OpenStudy (anonymous):

The price of a gallon of unleaded gas has risen to $2.89 today. Yesterdays price was 2.83. Find the percentage increase. Round to the nearest tenth of a percent

OpenStudy (anonymous):

0.06 =6

OpenStudy (hoblos):

2.12%

OpenStudy (anonymous):

hoblos is correct

OpenStudy (anonymous):

\[\frac{2.89-2.83}{2.83}*100=2.12\%\]

OpenStudy (anonymous):

0.06/before value = (x)*100% 2,12%